PWRDrEaMz:

Please help. Fine the missing number that makes the expression a perfect square. Next, write the expression in factored form. 64x^2-144x+___

PWRDrEaMz:

HALPPPP

AZ:

to make it a perfect square, it has to be in the form a^2 - 2ab + b^2 = (a-b)^2 we have 64x^2 - 144x + ____

AZ:

how can you rewrite 64x^2 so that way it's something ( )^2

PWRDrEaMz:

^2 means squared btw

AZ:

I'm glad you know that what squared gives you 64?

PWRDrEaMz:

8?

AZ:

Yes So do you agree that 64x^2 can be written as (8x)^2

PWRDrEaMz:

no not 8

PWRDrEaMz:

It's 8x

PWRDrEaMz:

cause that could make 64x^2

PWRDrEaMz:

Yes

AZ:

yes, 8x good so remember \(\color{red}{a}^2 - 2\color{red}{a}\color{green}{b} + \color{green}{b}^2 = (\color{red}{a} - \color{green}{b})^2\) and your equation can now be written as \( \color{red}{(8x)}^2 - 144x + \text{____}\) how do you find b?

AZ:

remember that 144x is equal to 2ab and we know that 'a' is 8x so what would be the 'b' part? 144x = 2*8x * b solve for b

PWRDrEaMz:

is b the ___

PWRDrEaMz:

missing number

AZ:

the blank is going to be b^2

AZ:

we're trying to solve for b^2 but first we have to figure out b

PWRDrEaMz:

Im trying to figure out the unfactored equation

PWRDrEaMz:

SO I need the FOIL equation to find the answer

PWRDrEaMz:

reeee

AZ:

yes, I know but first we have to figure out what 'b' is to complete the square to make it a perfect square, we need to find what 'b' is so then we add b^2 so that way it becomes a perfect square and can be simplified to (a-b)^2

AZ:

The FOIL equation will not help here

PWRDrEaMz:

oof

PWRDrEaMz:

88

PWRDrEaMz:

?

PWRDrEaMz:

nahh

PWRDrEaMz:

I always doubt myself :/

AZ:

no we know that the middle term is equal to 144x and that the middle term is equal to 2ab and a is 8x because a^2 gives us 64x^2 so we need to figure out what b is to make 2ab as 144x so plug in what we know 144x = 2*(8x) * b what is 2 * 8x =

PWRDrEaMz:

16x OOOOOOOOOOOOOHHHHH

PWRDrEaMz:

Tysm lmao sorry

AZ:

now divide that on both sides

AZ:

144x / 16x = ?

PWRDrEaMz:

oh wait

PWRDrEaMz:

9x

AZ:

we need to find 'b' so that way we can find b^2 and no, remember there's an x in the numerator and denominator so they both get cancelled out

PWRDrEaMz:

9

PWRDrEaMz:

I meant

PWRDrEaMz:

oh 81

AZ:

yes and so you're b is 9 so do you see how we can re-write it now? \(\color{red}{a}^2 - 2\color{red}{a}\color{green}{b} + \color{green}{b}^2 = (\color{red}{a} - \color{green}{b})^2\) \( \color{red}{(8x)}^2 - 2\color{red}{(8x)}\color{green}{(9)} + \text{____}\) we have to add it by b^2 to make a perfect square and we know that b = 9 so what is 9^2

PWRDrEaMz:

81

AZ:

there you go does what we did make more sense now?

PWRDrEaMz:

Yup

PWRDrEaMz:

kinda

AZ:

we had to find some number so that way it can be simplified into (a-b)^2 and so you take that middle number and divide it by 2a and then you square it

PWRDrEaMz:

Tysm tho

AZ:

No problem!

PWRDrEaMz:

And you helped me level up :D

PWRDrEaMz:

Im level 19 now :D

PWRDrEaMz:

ALMOST ORANGEEE

Quicksilver0125:

ur level 20

Quicksilver0125:

i need to at least get to level 10 plz help me

Quicksilver0125:

plz give me a medal